# RAG vs. Long Context: Examining Frontier Large Language Models for Environmental Review Document Comprehension

### v1

@@ 5 Related Work
Long Context Evaluation A popular technique to evaluate the long context LLMs is with a simple needle in a haystack analysis to test in-context retrieval ability Chandrayan et al. (2024). Despite the claims made in these tests, it is shown current LLMs perform poorly in processing and understanding long, context-rich sequences in rigorous scientific benchmarks Li et al. (2024); Liu et al. (2024). For example, Li et al. (2024) constructed LongICLBench benchmark to assess a set of long-context LLMs in an extreme-label classification task as a long in-context learning task. They showed that long context understanding and reasoning is still a challenging task for the existing LLMs. Few studies showed that long-context LLMs are affected by the position of the relevant information in the input context Liu et al. (2024); Ivgi et al. (2023). For instance, Ivgi et al. (2023) showed that encoder-decoder models have significantly higher performance when relevant information is placed at the start of the input context. In addition, Liu et al. (2024) showed that LLMs perform weakly when they must access relevant information in the middle of long contexts.
@@
RAG for Long Documents RAG models offer a promising approach for enabling LLMs to search and extract relevant information from lengthy documents or extensive collections. The common strategy of splitting documents into smaller, more manageable chunks that fit within the LLM’s context window has its limitations, as highlighted by recent studies  Barnett et al. (2024). These limitations include the model’s failure to accurately extract answers even when they are present within the provided context, particularly when there is excessive noise or contradictory information. To address these challenges, researchers have proposed novel techniques. HippoRAG, a neurobiologically inspired long-term memory system designed for LLMs to handle long documents more effectively, aims to mitigate the limitations of current RAG models  Gutiérrez et al. (2024).  Gao et al. (2023) provide a comprehensive survey of RAG methods, categorizing them into three paradigms: Naive RAG, Advanced RAG, and Modular RAG. The authors highlight the remarkable adaptability of Modular RAG, which allows for module substitution or reconfiguration to address specific challenges, surpassing the fixed structures of Naive and Advanced RAG. Modular RAG integrates new modules or adjusts interaction flow among existing ones, enhancing its applicability across different tasks. The survey also discusses the concept of adaptive retrieval in RAG, exemplified by methods like Flare  Jiang et al. (2023) and Self-RAG  Asai et al. (2023). These approaches refine the RAG framework by enabling LLMs to actively determine the optimal moments and content for retrieval, enhancing the efficiency and relevance of the sourced information. Despite these advancements,  Gao et al. (2023) emphasize that further research is needed to fully understand the intricacies of applying RAG to long documents and to develop more robust and reliable methods.
@@ 6 Conclusion
In this study, we conduct the initial investigation into the performance of LLMs within the domain of the National Environmental Policy Act and its associated documents. To facilitate this, we introduce NEPAQuAD, a question-and-answering benchmark designed to evaluate a model’s capability to understand the legal, technical, and compliance-related content found in NEPA documents. We assess three frontier LLMs designed for handling extensive contexts—Claude Sonnet, Gemini, and GPT-4—across various contextual settings. Our comprehensive analysis indicates that NEPA documents pose a significant challenge for LLMs, particularly in terms of understanding the complex semantics and effectively processing the lengthy documents. The findings reveal that models augmented with the RAG technique surpass those LLMs that are simply provided with the PDF content as long context. This suggests that incorporating more relevant knowledge retrieval processes can significantly enhance the performance of LLMs on complex document comprehension tasks like those found in the NEPA domain. In addition, we noticed that these LLMs struggle to use long input contexts to answer more difficult questions that require multiple steps of reasoning. For example, models performed best in answering closed questions and worst in answering divergent and problem-solving questions.
@@ 7 Limitations
Uncertainty of generated responses by LLMs. Due to budget constraints, we conducted only one phase of response generation across different configurations. This introduces a risk of uncertain outputs, meaning that LLMs might generate different responses each time, even with the same input, as demonstrated in another study Wagle et al. (2024). In future work, we plan to run LLMs multiple timee and analyze the effect of this uncertainty in response generation.

### v2

@@ 5 Related Work
RAG for Long Documents RAG models offer a promising approach for enabling LLM s to search and extract relevant information from lengthy documents or extensive collections. The common strategy of splitting documents into smaller, more manageable chunks that fit within the LLM’s context window has its limitations, as highlighted by recent studies  Barnett et al. (2024). These limitations include the model’s failure to accurately extract answers even when they are present within the provided context, particularly when there is excessive noise or contradictory information. To address these challenges, researchers have proposed novel techniques. HippoRAG, a neurobiologically inspired long-term memory system designed for LLM s to handle long documents more effectively, aims to mitigate the limitations of current RAG models  Gutiérrez et al. (2024).  Gao et al. (2023) provide a comprehensive survey of RAG methods, categorizing them into three paradigms: Naive RAG, Advanced RAG, and Modular RAG. The authors highlight the remarkable adaptability of Modular RAG, which allows for module substitution or reconfiguration to address specific challenges, surpassing the fixed structures of Naive and Advanced RAG. Modular RAG integrates new modules or adjusts interaction flow among existing ones, enhancing its applicability across different tasks. The survey also discusses the concept of adaptive retrieval in RAG, exemplified by methods like Flare  Jiang et al. (2023b) and Self-RAG Asai et al. (2023). These approaches refine the RAG framework by enabling LLM s to actively determine the optimal moments and content for retrieval, enhancing the efficiency and relevance of the sourced information. Despite these advancements,  Gao et al. (2023) emphasize that further research is needed to fully understand the intricacies of applying RAG to long documents and to develop more robust and reliable methods.
@@ 6 Conclusion
In this study, we conduct the initial investigation into the performance of LLM s within the NEPA domain and its associated documents. To facilitate this, we introduce NEPAQuAD, a question-and-answering benchmark designed to evaluate a model’s capability to understand the legal, technical, and compliance-related content found in NEPA documents. We assess five long-context LLM s (both closed- and open-source) designed for handling extensive contexts across various contextual settings.
Our comprehensive analysis indicates that NEPA documents pose a significant challenge for LLM s, particularly in terms of understanding the complex semantics and effectively processing the lengthy documents. The findings reveal that models augmented with the RAG technique surpass those LLM s that are simply provided with the PDF content as long context. This suggests that incorporating more relevant knowledge retrieval processes can significantly enhance the performance of LLM s on complex document comprehension tasks like those found in the NEPA domain. In addition, we noticed that these LLM s struggle to use long input contexts to answer more difficult questions that require multiple steps of reasoning. For example, models performed best in answering closed questions and worst in answering divergent and problem-solving questions. However, when these models are augmented with additional context, as seen in the RAG setup, their performance can be significantly improved. We encourage the research community to focus on advancing the complementary use of long-context LLM s and RAG techniques together, rather than treating them as distinct, separate approaches.
@@ 7 Limitations
Uncertainty of generated responses by LLMs. Due to budget constraints, we conducted only one phase of response generation across different configurations. This introduces a risk of uncertain outputs, meaning that LLMs might generate different responses each time, even with the same input, as demonstrated in another study Wagle et al. (2024). In future work, we plan to run LLMs multiple times and analyze the effect of this uncertainty in response generation.
@@
Random errors when using RAGAS We noticed that RAGAS evaluation framework Es et al. (2023) produces random errors and "NaN" values in specific types of question. While running experiments to analyze the details of this fall outside of scope of this work, our preliminary tests lead us to assume that it might be due to the way the framework interacts with generative models and tries to conform their output to its formats.